# ArtCrafter: Text-Image Aligning Style Transfer via Embedding Reframing

### v1

@@ 4.2 Qualitative Evaluations
4.3 Quantitative Evaluations
As shown in Table 1, our method outperformed other text-guided stylization methods in the CLIP-Text, CLIP-Image, DINO-v2, and LPIPS metrics. This indicates that our method has a significant advantage in generating images that are highly consistent with the target description and style. Moreover, our method effectively retains the details and semantics of the content while maintaining high stylization quality as well as diversity.
4.4 User Study
To obtain a more comprehensive assessment, we conducted a user study. We randomly selected $30$ generated results for each method covering a wide range of styles and hired $26$ professionals in the art field to evaluate these generated results. Specifically, they rated text and image consistency on a scale of $1-5$, resulting in the Human-Text and Human-Image scores in Table  1. The images with the best overall results are then pulled out to obtain the Human-Overall percentage results. The results in rows $6-8$ of Table 1 indicate that the results generated by our proposed ArtCrafter are more favored in all three aspects. We notice the difference between objective evaluation metrics and subjective evaluation metrics because the former assesses each aspect in isolation. Whereas users may integrate information across various aspects, despite separate options provided. Human Preference suggests that our generated results have struck a better balance among text consistency, image consistency, and overall visual appeal.
W.O. Sec. 3.2 23.49 66.81 37.25 0.4971
W.O. Sec. 3.3 19.13 70.52 41.77 0.4316
W.O. Sec. 3.4 22.19 71.39 42.63 0.3572
Table 2: Quantitative ablation study of proposed components.
Figure 5: Visual differences between ArtCrafter generated samples with global features and attention-based style features.
4.5 Ablation Study
@@
quantitative analysis. To fully evaluate the performance of the proposed components, we conduct an ablation study using both quantitative and qualitative methods. Table 2 summarizes the quantitative results, in which we test ArtCrafter’s performance after removing each essential technique component individually. Specifically, the metrics of CLIP-Text show that the attention-based style extraction component significantly enhances image consistency. Meanwhile, the evaluation results of CLIP-Image demonstrate that the text-image aligning augmentation component plays a positive role in guiding the text content. In addition, the results of measuring the similarity between the generated results using LPIPS indicate that the explicit modulation component plays a key role in enhancing the diversity of the generated results. Together, these findings demonstrate the significance and efficacy of the separate components in the entire strategy.
Visualization analysis. We first analyze the specific impact of attention-based style extraction (Sec. 3.2) on the results, as shown in Fig. 5. By utilizing this component, we effectively capture more detailed style features (e.g., line textures on walls and paper materials), thus achieving greater accuracy and richness in the style transfer process.
@@
To evaluate the effectiveness of text-image aligning augmentation (Sec. 3.3), we provide a visual representation of the stylization results in Fig. 6 and Fig. 7. Fig. 6 demonstrates that the text-image aligning augmentation component effectively contributes to the efficacy of the text role. In addition, the heatmap in Fig. 7 reveals the cosine similarity between the cross-attention of the generated results and the textual self-attention results, suggesting that ArtCrafter, through a fine-grained setting of the attention mechanism, makes the generated images more compatible with what the text describes. This visual evidence supports the success of our strategy for achieving tighter compatibility and refinement between text and images via text-image aligning augmentation module.
@@
Finally, we observe the effect of dynamic scale $\alpha$ in explicit modulation (Sec. 3.4) on the diversity of the generated results by adjusting it. As shown in Fig. 8, increasing its effect produces a more diverse set of images, but it still shares common stylistic attributes with the reference image. We typically set $\alpha$ to $0.8$.
Figure 9: Visualization results of ArtCrafter architecture trained under different datasets respectively.
Dataset. We construct the ArtMarket dataset using art images from WikiArt [7] and LAION-Aesthetics [21], and BLIP-2 [22] as the textual style descriptive model to build the art-description data pairs. We perform the ablation study by training the model on the LAION-Aesthetics dataset. Fig. 9 shows that the model trained only on LAION-Aesthetics is more inclined to generate realistic style landscape images, while our trained model effectively generates the style of Van Gogh’s Starry Night. This demonstrates that incorporating artistic data improves the model’s performance in capturing artistic styles.
4.6 Additional Analysis
Figure 10: The visual presentations of the pre-trained diffusion model influenced by ArtCrafter at varying intensities.
Figure 11: The combination of ArtCrafter and additional conditions. The results of this combination show that ArtCrafter can accept images as content guidance, achieve style transfer effects, and be compatible with 3D content information.
@@ 4.6 Additional Analysis
ArtCrafter control. We test the role of ArtCrafter in the pre-training diffusion model. The result of adjusting the scale of ArtCrafter enhanced embedding $E_{text{P}}$ in the pre-trained diffusion model is shown in Fig. 10. In this work, we usually set the scale to $0.6$ to balance content and style. This verifies that ArtCrafter can effectively enhance the applicability of pre-trained diffusion models in the art generation domain at a lower training cost. With this adjustment, we can more flexibly control the degree of text consistency and stylization of the generated images to meet different creative needs.
@@
ArtCrafter with additional conditions applied. Benefiting from our design without any changes to the network structure of the original diffusion model, ArtCrafter is seamlessly compatible with existing controllable tools [50]. Fig. 11 shows the diverse examples generated by applying different structural controls, including canny edge detection [3], normal map [40], HED edge detection [45], lineart edge, and MiDaS depth map [32]. These examples not only highlight ArtCrafter’s flexibility in adapting to a variety of conditions but also foretell its great potential for application in the $3D$ field, opening up new possibilities for artistic creation and visual design.
5 Conclusion
@@
Supplementary Material
A Additional Results
Fig. 1 illustrates our approach to transferring the “ink painting” style from a reference image to iconic global landmarks. For the references, users need not include terms like “painting,” “art,” “ink painting,” or other explicit artistic descriptors; instead, they can input straightforward content descriptors, such as “Neuschwanstein Castle, Bavaria.” This simplifies the user’s interaction process. Furthermore, Figs 910111213141516, and 17 display the style transfer effects on subjects including people, animals, and landscapes.
Figure 2: Additional qualitative comparison with other state-of-the-art text-guided stylization methods.
Figure 3: Quantitative comparison visualization.
B Additional Comparisons
Quanlitative comparison. In Fig. 2, we present an expanded set of comparative results with the current state-of-the-art methods, encompassing a diverse array of styles and textual inputs. These thorough comparative analyses underscore the robustness and superiority of ArtCrafter, showcasing its adept handling of varied styles and textual contents. The comparisons not only manifest ArtCrafter’s proficiency in executing style transfer efficiently but also accentuate its precision in preserving the integrity of text content.
Quantitative comparison visualization. In Fig. 3, we employ scatter plots to visualize the performance of the state-of-the-art methods in terms of the quantitative metrics CLIP-Image [31] and CLIP-Text. This visualization effectively demonstrates how our method strikes a balance between stylistic and content expression, thereby showcasing its exceptional performance. By plotting the results on these metrics, we provide a clear and concise comparison that illustrates the strengths of ArtCrafter in achieving a harmonious blend of style and content, which is a critical aspect in the field of style transfer generation.
Figure 4: Dataset samples. A collection of artworks and corresponding textual descriptions.
C ArtMarket Dataset
Our dataset, ArtMarket, depicted in Fig. 4, comprises a rich collection of artworks accompanied by their corresponding textual descriptions. The visual content of the dataset is primarily sourced from the extensive art repository of WikiArt [7] and LAION-Aesthetics [21], which features a diverse range of artistic styles and periods. The descriptive texts are derived from BLIP-2 [22], a sophisticated language model that provides detailed and contextually rich captions. This combination of high-quality images and descriptively accurate texts allows ArtMarket to effectively support the training of models aimed at style transfer and content generation in the realm of text-to-image synthesis.
Figure 5: Style transfer results with text-image aligning augmentation ablation study.
Figure 6: Diversity results with explicit modulation ablation study.
Figure 7: Ablation study of ArtCrafter guidance scale.
D Additional Ablation Study
Text-image aligning augmentation. Fig. 5 presents an array of additional ablation study results for our text-image aligning augmentation component, underscoring its role in bolstering the efficacy of text-guided content generation. These results demonstrate that our component significantly enhances the model’s ability to produce content that aligns closely with the textual prompts, without compromising the stylistic integrity of the image. The experiments showcase how the integration of our aligning augmentation leads to a harmonious fusion of content fidelity and style preservation.
Diversity. Fig. 6 elaborates on the ablation study results for our Explicit Modulation component, further substantiating its effectiveness in augmenting the diversity of content generation. These results highlight the component’s capacity to enrich the variety of outputs produced by the model, ensuring a broader spectrum of creative expressions. The experiments detailed in Fig. 6 underscore how explicit modulation contributes to the generation of images that not only capture the essence of the input text but also exhibit a wide range of artistic styles, thus validating its importance in achieving diverse and compelling visual outcomes.
Guidance scale. We conduct comprehensive validation of the integration of ArtCrafter composite modules into the U-Net structure of the diffusion model. This validation serves as a supplementary extension to the main experimental results presented in the text, offering an in-depth analysis of the model’s performance across a range of guidance intensities. Specifically, Fig. 7 illustrates the outcomes at intervals of $0.1$ for the guidance coefficient, thereby providing a granular view of how varying levels of input influence the final generation quality. These additional results underscore the robustness and flexibility of ArtCrafter approach in adapting to different degrees of user-specified control.
Figure 8: Failure cases. Instances of ineffective style transfer.
E Failure Cases
Our work exhibits a degree of instability in consistently capturing the precise locations of distinctive stylistic textures, as depicted in Fig. 8. The reference images contain extensive, curved textural elements, exemplified by the blue and pink curves, which pose a significant challenge for our model to fully assimilate and reproduce in the generated outputs. This observation highlights an area for future improvement, where enhancing the model’s ability to accurately replicate intricate, long-range textural patterns could lead to more stable and faithful style transfers. Despite this, the overall effectiveness of our approach in handling complex style elements is evident.
Figure 9: Additional results. Further demonstrations of style transfer efficacy.
Figure 10: Additional results. Further demonstrations of style transfer efficacy.
Figure 11: Additional results. Further demonstrations of style transfer efficacy.
Figure 12: Additional results. Further demonstrations of style transfer efficacy.
Figure 13: Additional results. Further demonstrations of style transfer efficacy.
Figure 14: Additional results. Further demonstrations of style transfer efficacy.
Figure 15: Additional results. Further demonstrations of style transfer efficacy.
Figure 16: Additional results. Further demonstrations of style transfer efficacy.
Figure 17: Additional results. Further demonstrations of style transfer efficacy.

### v2

@@ IV-B Qualitative Evaluations
Figure 5: Ablation studies for attention-based Style Extraction. (a) Visual differences between ArtCrafter generated samples without preceiver attention and our method. (b) Discussion on the number of learnable queries N.
@@
IV-C Quantitative Evaluations
As shown in Table I, our method outperformed other text-guided stylization methods in the CLIP-Text, CLIP-Image, DINO-v2, and LPIPS metrics. This indicates that our method has a significant advantage in generating images that are highly consistent with the target description and style. Moreover, our method effectively retains the details and semantics of the content while maintaining high stylization quality as well as diversity.
IV-D User Study
To obtain a more comprehensive assessment, we conducted a user study. We randomly selected $30$ generated results for each method covering a wide range of styles and hired $26$ professionals in the art field to evaluate these generated results. Specifically, they rated text and image consistency on a scale of $1-5$, resulting in the Human-Text and Human-Image scores in Table  II. The images with the best overall results are then pulled out to obtain the Human-Overall percentage results. The results of Table II indicate that the results generated by our proposed ArtCrafter are more favored in all three aspects. We notice the difference between objective evaluation metrics and subjective evaluation metrics because the former assesses each aspect in isolation. Whereas users may integrate information across various aspects, despite separate options provided. Human Preference suggests that our generated results have struck a better balance among text consistency, image consistency, and overall visual appeal.
IV-E Ablation Study
@@
Figure 9: Diversity results with explicit modulation ablation study. We conducted an ablation study to assess the impact of the explicit modulation component on the diversity of generated images. The results demonstrate that the explicit modulation component significantly enhances the diversity of the outputs.
IV-E1 Visualization analysis.
Attention-based style extraction. In our study, we observed that naive cross-attention tends to cause a shift in style, as shown in the results of Fig. 4 (Styleshot, Stylealigned, Stylecrafter, CSGO). The novel approach is by processing the embedded image’s latent features with perceptual attention (Sec. III-B). This method not only effectively prevents unintended style shifts but also significantly enhances artistic expression, particularly in terms of color and texture. As demonstrated in Fig. 5 (a), the images processed with perceptual attention show a remarkable improvement in the richness of color and the detail of texture, making the generated images more artistically appealing and visually attractive.
To further optimize the performance of our model, we conducted an in-depth ablation study on the key parameter $N$. As shown in Figure 5 (b), we found that the value of $N$ has a crucial impact on the model’s attention distribution and generation results. When $N$ is low, the model’s attention becomes overly focused on narrow feature regions, leading to the loss of feature information and resulting in images that appear overly localized and lack a sense of coherence. Conversely, when $N$ is high, the model’s attention becomes too dispersed, blurring the key attributes of the noise and causing the generated images to lack clear structure and detail. Through this series of experiments, we identified an optimal balance point, enabling the model to capture and reproduce key image features accurately while maintaining the overall style.
TABLE III: Quantitative ablation study of proposed components. The results provide insights into how each component affects the performance of key metrics such as image consistency, text consistency, and diversity.
W.O. ASE III-B 23.49 66.81 37.25 0.4971
W.O. TIAA III-C 19.13 70.52 41.77 0.4316
W.O. EM III-D 22.19 71.39 42.63 0.3572
@@
Text-image aligning augmentation. To evaluate the effectiveness of the TIAA (Sec. III-C), we present visual representations of the stylization results in Fig. 6 and Fig. 7. Fig. 6 clearly illustrates that the TIAA component significantly enhances the alignment between text and image, thereby improving the overall efficacy of the text role. Moreover, the heatmap in Fig. 7 reveals the cosine similarity between the cross-attention of the generated results and the textual self-attention results. This indicates that ArtCrafter, through its fine-grained attention mechanism, effectively aligns the generated images with the textual descriptions. This visual evidence strongly supports the success of our strategy in achieving tighter compatibility and refinement between text and images via the TIAA module.
Figure 10: Visual demonstrations of ArtCrafter’s influence on pre-trained diffusion models. The illustrations reveal how ArtCrafter’s modulation enhances the model’s ability to generate images that are stylistically coherent with varying degrees of influence. As the intensity of ArtCrafter’s application increases, there is a noticeable evolution in the image’s stylistic attributes, demonstrating the adaptability and controllability of our method.
Figure 11: Visualization results of ArtCrafter architecture trained on different datasets. Specifically, when trained on a dataset with rich artistic styles, ArtCrafter generates images with intricate and varied artistic features. In contrast, when trained on a dataset with more natural images, the model produces outputs that are highly consistent with natural visual elements. These results underscore the robustness and versatility of ArtCrafter in generating high-quality images.
@@
Explicit modulation. We investigate the impact of the dynamic scale $\alpha$ in EM (Sec. III-D) on the diversity of generated results by adjusting its value. As illustrated in Fig. 8, increasing $\alpha$ leads to a more diverse set of images while maintaining common stylistic attributes with the reference image. We typically set $\alpha$ to $0.8$.
Fig. 9 provides detailed ablation results for our EM component, further demonstrating its effectiveness in enhancing content diversity. These results highlight the component’s ability to significantly enrich the variety of outputs, ensuring a broader spectrum of creative expressions. The experiments detailed in Fig. 9 underscore how EM contributes to generating images that not only capture the essence of the input text but also exhibit a wide range of artistic styles, thereby validating its importance in achieving diverse and compelling visual outcomes.
IV-E2 quantitative analysis.
To fully evaluate the performance of the proposed components, we conduct an ablation study using both quantitative and qualitative methods. Table III summarizes the quantitative results, in which we test ArtCrafter’s performance after removing each essential technique component individually. There are inherent trade-offs between image consistency, text consistency, and diversity. Each module enhances a specific aspect: ASE improves image performance (CLIP-Image, DINO-v2), TAA boosts text performance (CLIP-Text), and EM parametrically balances and diversifies the results (LPIPS). Our goal is to achieve an optimal balance among these aspects to maximize overall effectiveness. Collectively, these findings underscore the significance and efficacy of each component within the overall strategy.
IV-E3 ArtCrafter control.
@@
We test the role of ArtCrafter in the pre-training diffusion model. The result of adjusting the scale $\alpha$ of ArtCrafter in the pre-trained diffusion model is shown in Fig. 10. Fig. 10 illustrates the outcomes at intervals of $0.1$ for the guidance coefficient, thereby providing a granular view of how varying levels of input influence the final generation quality. In this work, we usually set the scale to $0.6$. This verifies that ArtCrafter can effectively enhance the applicability of pre-trained diffusion models in the art generation domain at a lower training cost. With this adjustment, we can more flexibly control the degree of text consistency and stylization of the generated images to meet different creative needs.
IV-E4 Dataset.
We constructed the ArtMarket dataset using art images from WikiArt [49] and LAION-Aesthetics [50], and utilized BLIP-2 [46] as the textual style descriptive model to create art-description data pairs. To evaluate the impact of our dataset, we conducted an ablation study by training the model solely on the LAION-Aesthetics dataset. As shown in Fig. 16, the model trained only on LAION-Aesthetics tends to generate realistic landscape images. In contrast, our model trained on the combined dataset effectively captures the distinctive style of Van Gogh’s Starry Night. This demonstrates that incorporating diverse artistic data significantly enhances the model’s ability to generate a wide range of artistic styles.
Figure 12: Integration of ArtCrafter with additional conditions. The outcomes of this integration indicate that ArtCrafter not only effectively incorporates images as content guidance for style transfer but also seamlessly integrates with 3D content information. This compatibility showcases ArtCrafter’s potential to enhance the richness of generated images by leveraging various types of input data, thereby expanding its applicability in diverse creative scenarios.
V Discussion
V-A Application
V-A1 ArtCrafter with additional conditions applied.
@@ V-A Application
Benefiting from our design without any changes to the network structure of the original diffusion model, ArtCrafter is seamlessly compatible with existing controllable tools [25]. Fig. 12 shows the diverse examples generated by applying different structural controls, including canny edge detection [51], normal map [52], HED edge detection [53], lineart edge, and MiDaS depth map [54]. These examples not only highlight ArtCrafter’s flexibility in adapting to a variety of conditions but also foretell its great potential for application in the $3D$ field, opening up new possibilities for artistic creation and visual design.
Figure 13: Enhanced comparisons with IP-Adapter variants. To further validate the superiority of our approach, we conducted additional comparative analyses with the SD1.5 and Flux versions of the IP-Adapters, which were enhanced with our TIAA and EM algorithms. These comparisons provide a comprehensive view of how our algorithms contribute to the performance of these models.
V-A2 More applications on other T2I models.
To further verify the transferability and universality of the proposed method in multi-functional application scenarios, we conducted a series of in-depth comparative experiments on various text-to-image generation models with different weight configurations and architectural designs. As shown in Fig. 13, we present the comparative results of two representative model versions—Stable Diffusion 1.5 (SD1.5) and Flux—after integrating two key modules of our method: the TIAA module and the EM module. The experimental results demonstrate that by embedding these two modules into the models in a lightweight integration manner, significant performance improvements can be achieved without the need for additional training or fine-tuning. This fully proves the transferability and efficiency of our method.
Figure 14: Versatility of our algorithms in natural image editing. Our algorithms are designed to be versatile and applicable across a broad spectrum of natural image editing applications. They are effective regardless of the artistic dataset used, demonstrating their robustness and adaptability. This capability allows for a seamless integration into various workflows, enhancing the potential for creative expression.
V-A3 Natural image editing.
The strengths of ArtCrafter are not confined to its training on the ArtMarket dataset; its performance in various natural image editing scenarios is also highly impressive. Fig. 14 demonstrates the remarkable capabilities of TIAA and EM in jointly enhancing image-text guidance. Our method can be widely applied to multiple natural image editing scenarios, including scene editing (such as adjusting the atmosphere or weather of a scene), adding objects (naturally integrating new objects into an image), non-rigid editing (deforming objects), and object replacement (replacing one object in an image with another). In Fig. 14 (a), the improved results more accurately reflect the ”road” information from the text, such as more clearly presenting the texture, direction, or surrounding environment of the road. In Fig. 14 (b), (c), and (d), not only are the requirements of the editing text better enhanced, but the information of non-edited parts, such as the details of the indoor environment, is also preserved, ensuring the naturalness and coherence of the overall image. This capability makes ArtCrafter valuable and broadly applicable in the field of natural image editing.
Figure 15: Enhanced comprehension of extended text. Our model demonstrates an advanced ability to interpret and understand lengthy and intricate textual descriptions.
Figure 16: Dataset samples. The dataset samples cover a diverse range of artistic styles and periods, ensuring a comprehensive representation of the artistic spectrum. The textual descriptions are crafted to capture the essence of each artwork, providing detailed insights into the visual elements, thematic content, and stylistic nuances.
V-A4 Long Text Prompt
In Fig. 15, we present a series of long and complex textual requirements that our model successfully translates into detailed visual outputs. For instance, in Fig. 15 (a), the model accurately generates an image with a deer having a “curious expression with big eyes and a wet nose just above the waterline”, “detailed fur texture”, “and ripples in the water”, capturing the “half above and half below water surface” scenario. In Fig. 15 (b), the model creates a serene forest scene with a deer “surrounded by lush foliage,” “a river meandering,” and “clear water reflecting the sky.” These examples demonstrate the robust capability of our text-image alignment design in handling intricate textual instructions and generating images that precisely match the detailed descriptions.
Figure 17: Analysis of failure cases. (a) Text generation failed. (b) Instances of inconsistent line texture.
V-B Dataset Samples
Our dataset ArtMarket, as shown in Fig. 16, contains a rich collection of artworks along with their corresponding textual descriptions. The visual content of the dataset is primarily sourced from the extensive art repositories of WikiArt [49] and LAION-Aesthetics [50], which cover a wide range of artistic styles and periods. The descriptive texts are derived from BLIP-2 [46], an advanced language model that can generate detailed and contextually rich captions. The combination of high-quality images and accurately descriptive texts enables ArtMarket to effectively support the training of models for style transfer and content generation in the field of text-to-image synthesis.
V-C Limitation
The limitations of ArtCrafter are mainly reflected in two aspects. First, it is the ability to accurately generate text. As shown in Fig. 17 (a), ArtCrafter has difficulty in accurately generating text in images and cannot ensure the clarity and accuracy of the text content “ArtCrafter”. This limitation may seriously affect its performance in image generation tasks that require text elements. For example, in designing posters, brochures, or other image editing tasks that require precise text layout, ArtCrafter may not be able to meet the high-precision requirements of users for text content and layout. Secondly, the model can only partially follow large-scale textures and brushstrokes. For example, in Fig. 17 (b), our method incorrectly transforms the large-scale curves from left to right in the image into inconsistent curve shapes. This error indicates that the model may fail to fully understand and accurately reproduce the original shapes and directions when dealing with complex textures and brushstrokes. Accurately replicating complex long-range texture patterns is one of the challenges for our model to fully assimilate and reproduce in the generated output.
VI Conclusions and Future Work